OpenStudy (anonymous):

Which is the simplified form of 15 times q times r to the fifth power all over 3 times q cubed times r to the fourth power.? @Nurali

OpenStudy (kainui):

Try to write it out as an equation. Slowly transfer everything from words into symbols as you read it. I'll help you along, don't be afraid about being wrong.

OpenStudy (anonymous):

15qr^5/3q^3r^4

OpenStudy (kainui):

Looks perfect to me, so now you need to simplify it. Remember that if you ever have trouble with exponents that they really just mean it's that number of them multiplied together. So for example: \[\frac{ x^5 }{ x^2 }=\frac{ x*x*x*x*x }{ x*x }=x*x*x=x^3\] So now you can see why you subtract the bottom exponents from the top exponents. They just divide out like 1's! Also, you can see that subtracting is also where x^(-1) = 1/x comes from.

OpenStudy (anonymous):

I don't understand

OpenStudy (kainui):

What don't you understand?

OpenStudy (anonymous):

That last part..

OpenStudy (nurali):

\[\frac{ 15qr^5 }{ 3q^3r^4 }\] \[\frac{ 5qr^5 }{ q^3r^4 }\] \[5q ^{1-3}r ^{5-4}\] \[5q ^{-2}r ^{1}\] \[\frac{ 5r }{ q^2 }\]